# Natural Language Dataset Generation Framework for Visualizations Powered by Large Language Models

### v1

@@ Abstract.
We introduce a Large Language Model (LLM) framework that generates rich and diverse NL datasets using only Vega-Lite specifications as input, thereby streamlining the development of Natural Language Interfaces (NLIs) for data visualization. We propose two techniques to synthesize relevant chart semantics accurately and enhance syntactic diversity in each NL dataset, respectively: 1) a guided discovery incorporated into prompting so that LLMs can steer themselves to create varying NL datasets in a self-directed manner; 2) a score-based paraphrasing to augment NL syntax along with four well-defined language axes. We also present a new chart collection of 1,981 real-world Vega-Lite specifications that have increased diversity and complexity compared to benchmarks, to demonstrate the generalizability of our framework. The experimental results show that our framework accurately extracts chart semantics and generates L1/L2 captions with 89.4% and 76.0% accuracy, respectively, while generating and paraphrasing utterances and questions with greater diversity than benchmarks. The codes and chart collection are available at https://github.com/hyungkwonko/chart-llm.
@@ 1. Introduction
Recent advancements in Natural Language Processing (NLP) techniques empowered individuals with limited data analysis and visualization expertise to engage in text-based interaction and execute data visualization tasks (Sultanum and Srinivasan, 2023; Shen et al., 2023). Because NL interaction methods are recognized as more natural and user-friendly compared to traditional approaches like direct manipulation (Setlur et al., 2016), many studies have incorporated Natural Language Interfaces (NLIs) into their systems. For example, users can query key data insights within charts using NL sentences (Srinivasan et al., 2018), significantly reducing the need for manual extraction via programming. Furthermore, they can input text to receive automatic recommendations for the most appropriate chart types (Narechania et al., 2020; Dibia, 2023), rather than selecting effective representations manually based on graphical language criteria (Mackinlay, 1986).
The problem is that numerous publications have consistently highlighted the scarcity of high-quality chart collections and NL datasets (Chen and Liu, 2023; Shen et al., 2022; Davila et al., 2020; Hutchinson et al., 2023; Luo et al., 2021a; Srinivasan et al., 2021) to develop NLIs for visualizations. Among many factors that explain the quality of NL datasets, securing syntactic diversity is pivotal (Gao et al., 2015; Setlur et al., 2016; Yu and Silva, 2019) to capture the language variations arising from diverse spectrum of user expertise, usage scenario, and personal preferences. However, it is challenging to create NL datasets designed for specific tasks with varying syntax through crowdsourcing, as excluding inattentive participants may diminish sample diversity, and overlooking inattention can introduce spurious outcomes alongside heightened statistical noise (Thomas and Clifford, 2017). Moreover, to annotate diverse and complex charts, researchers must disassemble the task into discrete subtasks to attain a faithful dataset (Kittur et al., 2011). Considering the distinct characteristics of each NL dataset (Shen et al., 2022), they need to devise multiple crowdsourcing initiatives, which entails additional efforts that scale proportionally with the number of dataset types. Similarly, it is noted only a fraction of chart collections (17 out of 56) is publicly accessible. While most of them exhibit limited diversity (e.g., in terms of chart type) (Chen and Liu, 2023), it is also a crucial factor in generating a variety of NL datasets, since many of them are derived from existing chart collections (Lundgard and Satyanarayan, 2021; Srinivasan et al., 2021; Kim et al., 2020).
We introduce an LLM framework for generating various NL datasets from Vega-Lite specifications (Satyanarayan et al., 2017) (Figure 2). Our framework comprises of two novel prompting techniques to synthesize relevant chart semantics accurately and to enhance syntactic diversity, respectively. First, we leverage guided discovery (Brown and Campione, 1994) so that LLMs can steer themselves to create varying NL datasets in a self-directed manner. In detail, the framework analyzes and integrates chart semantics (e.g., mark, encoding) with our scaffolding in accordance with the characteristics of each NL dataset. Also, by answering on key questions, it autonomously concentrates on the chart’s key features or propose high-level decisions. Second, we utilize score-based paraphrasing (Table 3) to increase the syntactic diversity of the generated NL datasets. We define four language axes (clarity, expertise, formality, subjectivity) using automatic qualitative coding (Hämäläinen et al., 2023), and the framework augments the generated datasets quantitatively along these axes.
@@
We also present a new collection of 1,981 Vega-Lite specifications (Figure 3), which is used to demonstrate the generalizability and viability of our NL generation framework. This collection is the largest set of human-generated charts obtained from GitHub to date. It covers varying levels of complexity from a simple line chart without any interaction (i.e., simple charts) to a chart with four plots where data points are linked with selection interactions (i.e., extra complex charts) (see the charts highlighted with red stroke in Figure 3). As we focus on amassing a richer set of charts in terms of complexity, more than 86% of them are in complex and extra complex levels. Compared to the benchmarks, our dataset shows the highest average pairwise edit distance between specifications, which proves that the charts are highly diverse from one another. Moreover, it contains the largest number of charts with composite views, interactions (e.g., tooltips, panning & zooming, and linking), and diverse chart types (e.g., map, grid & matrix, diagram, etc.) (Table 5).
Through our experiment, we show that using Vega-Lite specifications in conjunction with our prompting techniques for LLMs yields NL datasets of high fidelity. Specifically, we have generated L1/L2 captions (Lundgard and Satyanarayan, 2021), utterances with different phrasings (command, query, question) (Srinivasan et al., 2021), questions with varying types (lookup/compositional, visual/non-visual, open-ended) (Kim et al., 2020; Hoque et al., 2022). We assessed the accuracies of both the chart semantics and L1/L2 captions. We also measured the within- and cross-distribution metrics of utterances and questions compared to the benchmark datasets (Srinivasan et al., 2021; Kim et al., 2020). Our experiments showed that the accuracy of the analyzed chart semantics and generated L1/L2 captions is 89.4% and 76.0%, respectively. For the semantic diversity, the generated and paraphrased NL datasets had greater diversity in terms of 4.75 out of 6 within-distribution metrics on average.
@@
  • We present an LLM framework to generate various NL datasets leveraging guided discovery to synthesize relevant chart semantics accurately;
  • We introduce a score-based paraphrasing technique to enhance the syntactic diversity of NL datasets;
  • We propose a collection of 1,981 real-world Vega-Lite specifications that are highly diverse and complex compared to the benchmarks.
@@ 2. Background and Related Work
In this section, we explain the types of NL datasets that are of particular interest in the context of this work. We also explain the Vega-Lite specification and use of LLMs in synthesizing NL datasets.
@@ 2.3. LLMs and NL Datasets
3. NL Generation Framework
Figure 2. LLM Framework to Generate Varying NL Datasets for Visualizations. We start by minifying Vega-Lite specifications (b). Subsequently, we employ scaffolding and key questions (c), to generate L1/L2 captions, utterances, and questions (d). This is followed by score-based paraphrasing (e), allowing us to produce syntactically paraphrased NL datasets (f).
Table 1. Prompting techniques to generate each NL dataset. Each prompt is designed by choosing the most appropriate techniques considering the characteristics of different NL datasets.
@@ 3. NL Generation Framework
The goal of our framework is to generate high-quality NL datasets using only Vega-Lite specifications and prompt engineering. The framework consists of two main ideas (Figure 2); the first focuses on identifying relevant and accurate information through guided-discovery, while the second focuses on increasing syntactic diversity using score-based paraphrasing. To generate each type of NL dataset, we design each prompt to be maximally helpful by selecting the most appropriate strategies (Table 1).
@@ 3.2. Ensuring Accuracy and Relevance
Our framework leverages the concept of guided discovery (Brown and Campione, 1994) based on Chain-of-Thought prompting (Wei et al., 2022) to harness the maximum reasoning capability of LLMs. We employ two strategies of guided discovery: providing scaffolds (Hmelo-Silver et al., 2007). and posing key questions (De Jong and Van Joolingen, 1998). To analyze and integrate the chart semantics necessary for generating a specific NL dataset, we assist LLMs by offering scaffolds. Additionally, we furnish LLMs with key questions to guide their self-directed progress. This maximizes the use of LLMs’ reasoning abilities, allowing them to make decisions on which aspects to focus on and delve into when creating a particular data. Below, we denote each step where we italicize the relevant phrases, and utilize symbols for (#S) scaffolding and (#K) key question to make them easily identifiable.
@@ 3.2.1. L1 Caption
Considering real-world Vega-Lite specifications, we first understand whether the given chart is (#S) a composite view (e.g., layered, trellis, and multiple views), to enable top-down analysis of each chart one by one. The prompt follows a template with three questions to answer: Is it a composite view?; If it is, identify its type among layered, trellis, and multiple views; and determine the number of plots in the chart. Next, it analyzes each chart individually based on the provided scaffold of (#S) chart semantics: Data, Transform, Mark, Chart-Type, Encoding, Style, and Interaction, using the information about composite view. Here, we access the backing dataset to provide the (#S) fields that are presented in the Vega-Lite specification, along with their synonyms (i.e., (#S) titles also found in the same Vega-Lite specification) and their (#S) unique values if they are categorical variables. After analyzing all of these semantics, the LLM finally generates the L1 caption by combining them.
@@ 3.2.2. L2 Caption
L2 captions, unlike L1 captions that provide an overall description of the chart, offer the flexibility to selectively focus on specific features that capture the viewer’s interest. To craft informative and insightful captions, we follow a structured approach centered around a key question: (#K) What is the most prominent and meaningful feature in the given chart? Once we identify this feature, we delve deeper by exploring the mathematical operations required to analyze it: (#K) What is the mathematical operation(s) required to describe the feature? Subsequently, based on these operations, we generate (#K) a series of questions to analyze the feature (e.g., for the simple line chart with a red border in Figure 3, the following questions are generated: What was the highest stock price of Google?; What was the lowest stock price of Google?; What is the difference between the highest and lowest stock prices of Google?). This process allows us to create captions that provide valuable insights into the chart’s content. When answering questions, we utilize LangChain (Chase, 2022) to ensure that our captions contain only factual information in the captions. This step is crucial, as large language models (LLMs) have been known to produce hallucinations in response to mathematical problems (Ji et al., 2023). Once each question is answered, the collected information is subsequently incorporated into the final prompting stage for generating L2 captions. It’s important to note that, unlike previous work (Tang et al., 2023), we do not use any of the L1 captions when generating L2 captions. Instead, this is performed as an independent process.
@@ 3.2.3. Utterance
Similar to L1 captions, we begin by analyzing whether the chart is a (#S) composite view. We then proceed to generate (#S) instructions for each plot independently. This process entails creating a comprehensive set of step-by-step instructions for constructing each plot. To enhance readability and user-friendliness, we ensure that each instruction focuses on a single specific action, aligning with the same semantics used when generating L1 captions. For example, in the case of the simple line chart with a red border shown in Figure 3, the following instructions are generated:
@@
However, it’s important to note that the generated instructions may sometimes feature overly technical variable names from the chart, which might not align with users’ NL usage patterns. In such cases, we leverage information from synonyms found in the (#S) title of the Vega-Lite specification or (#S) values from the fields to replace these technical terms, resulting in more user-friendly instructions.
Next, we ask a key question to (#K) identify primary and secondary information. In this context, we anticipate that LLMs are able to automatically prioritize crucial semantics to paint a comprehensive picture of the chart, such as chart type or encoding, over additional instructions like style or interaction. This thought is based on Wang et al.’s observations (Wang et al., 2022), who noted that the typical workflow for creating visualizations often starts with this information (e.g., ’show me the price over time as a line chart’). Once we have all these components ready, we proceed to generate each type of utterance one by one, adhering to (#S) specific rules for each type. For commands, we employ the imperative voice. For queries, we use only variables, fields, attributes, mathematical formulas, abbreviations, and prepositions, while avoiding verbs and articles. For questions, we formulate inquiries in the form of questions. Across all types, we maintain the following rules: express each utterance in a single sentence, utilize only primary information, and keep the language concise and straightforward.

### v3

@@ Abstract.
We introduce VL2NL, a Large Language Model (LLM) framework that generates rich and diverse NL datasets using only Vega-Lite specifications as input, thereby streamlining the development of Natural Language Interfaces (NLIs) for data visualization. To synthesize relevant chart semantics accurately and enhance syntactic diversity in each NL dataset, we leverage 1) a guided discovery incorporated into prompting so that LLMs can steer themselves to create faithful NL datasets in a self-directed manner; 2) a score-based paraphrasing to augment NL syntax along with four language axes. We also present a new collection of 1,981 real-world Vega-Lite specifications that have increased diversity and complexity than existing chart collections. When tested on our chart collection, VL2NL extracted chart semantics and generated L1/L2 captions with 89.4% and 76.0% accuracy, respectively. It also demonstrated generating and paraphrasing utterances and questions with greater diversity compared to the benchmarks. Last, we discuss how our NL datasets and framework can be utilized in real-world scenarios. The codes and chart collection are available at https://github.com/hyungkwonko/chart-llm.
@@ 1. Introduction
Recent advancements in Natural Language Processing (NLP) techniques empowered individuals with limited data analysis and visualization expertise to engage in text-based interaction and execute data visualization tasks (Sultanum and Srinivasan, 2023; Shen et al., 2023). Many studies have incorporated Natural Language Interfaces (NLIs) into their systems to augment more natural and user-friendly interactions (Setlur et al., 2016). For example, Voder (Srinivasan et al., 2018) enables querying key data insights within charts using NL sentences, significantly decreasing the reliance on manual programming for data retrieval. Furthermore, users can provide text to receive automatic recommendations for the most appropriate chart types (Narechania et al., 2020; Dibia, 2023), rather than selecting effective representations manually based on graphical language criteria.
While the presence of suitable datasets modeling human behaviors is crucial in developing effective NLIs or tools for visualizations, prior work has repeatedly pointed to the scarcity of sizable pairs of high-quality datasets (chart, NL) (Chen and Liu, 2023; Shen et al., 2022; Davila et al., 2020; Hutchinson et al., 2023; Luo et al., 2021a; Srinivasan et al., 2021). In detail, existing chart datasets are occasionally synthetic (Zhao et al., 2020; Luo et al., 2021a), limited in diversity (e.g., chart type) (Mahinpei et al., 2022; Chen and Liu, 2023), or are limited to simpler charts (e.g., basic bar charts, univariate line charts) (Dibia and Demiralp, 2019). Making things worse, only a fraction of chart collections (17 out of 56) is publicly accessible (Chen and Liu, 2023). Furthermore, prior work collects the NL datasets that goes with the visualizations through crowdsourcing (Srinivasan et al., 2021). However, the process can be costly and time-consuming as it requires recruiting specific sets of target users of the system, some of whom must meet notably stringent qualification criteria. Moreover, it is challenging to capture the language variations that arise from a diverse spectrum of user expertise, usage scenarios, and personal preferences, although this is essential for addressing the syntactic variations among the target users of the systems in the real-world (Gao et al., 2015; Setlur et al., 2016; Yu and Silva, 2019). What exacerbates the situation is there are multiple types of NL tasks (e.g., captioning, chart generation & modification, and chart question-answering) where each one necessitates a new dataset tailored to the specific task or transfer knowledge.
@@
We present a new collection of 1,981 Vega-Lite specifications (Figure 2). This is the largest set of human-generated charts obtained from GitHub to date. It covers varying levels of complexity from a simple line chart without any interaction (i.e., simple charts) to a chart with four plots where data points are linked with selection interactions (i.e., extra complex charts) (see the charts highlighted with red stroke in Figure 2). As we focus on amassing a richer set of charts in terms of complexity, more than 86% of them are in complex and extra complex levels. Compared to the benchmarks, our dataset shows the highest average pairwise edit distance between specifications, which proves that the charts are highly diverse from one another. Moreover, it contains the largest number of charts with composite views, interactions (e.g., tooltips, panning & zooming, and linking), and diverse chart types (e.g., map, grid & matrix, diagram, etc.) (Table 2).
We also introduce VL2NL, a 3-stage NL generation framework that can be generalized to various NL tasks on visualizations (Figure 3). First, the framework preprocesses the underlying datasets and minifies Vega-Lite specification for efficient and effective usage by an LLM. Next, the framework leverages guided discovery (Brown and Campione, 1994) so that LLMs can steer themselves to create varying NL datasets in a self-directed manner. Here, it analyzes and integrates chart semantics (e.g., mark, encoding) with our scaffolding in accordance with the characteristics of each NL dataset. Also, by answering on key questions, it autonomously concentrates on the chart’s key features or propose high-level decisions. Finally, the framework applies a score-based paraphrasing (Table 5) with an LLM to simulate and include syntactic variations of human language in NL datasets.
To test VL2NL, we generated L1 captions that simply describe how the chart encodes data, L2 captions that describe the statistical properties of the data in a chart (Lundgard and Satyanarayan, 2021), utterances for chart generation (Srinivasan et al., 2021), and questions for chart question answering (Kim et al., 2020; Hoque et al., 2022). Our experiments showed that the accuracy of the analyzed chart semantics and generated L1/L2 captions is 89.4% and 76.0%, respectively. Moreover, the generated and paraphrased NL datasets showed greater syntactic diversity in terms of 4.75 out of 6 within-distribution metrics on average. Last, we demonstrate the application of our NL datasets in finetuning experiments, and the use of VL2NL in both fully-automatic and mixed-initiative modes within an interactive system for real-world scenarios.
@@
  • We collect 1,981 real-world Vega-Lite specifications that are diverse and go beyond simple charts;
  • We present 3-stage NL dataset generation framework for visualizations powered by LLMs that employs guided discovery and score-based paraphrasing;
  • We perform quantitative and qualitative analysis on the NL datasets generated by our framework.
@@ 2. Background and Related Work
In this section, we explain Vega-Lite specification and existing chart collections. Next, we present the types of NL datasets that are of particular interest in the context of this work. Last, we explain the use of LLMs in synthesizing NL datasets.
2.1. Chart Datasets
@@ 2.3. LLMs and NL Datasets
3. Vega-Lite Dataset
# of URLs / Vega-Lite specs
(a) URLs crawled 67,789
(b) URLs w/o duplicate 18,420
(c) URLs w/ license 7,408
(d) Specs after manual inspection 1,981
Quantity # of specs 4,318 9,897 6,680 52 709 1,981
Average # of keys in a spec 24 15 15 15 37 54
Simple (key $\leq$ 16) 0 6,164 6,354 41 186 73
Medium (key $\leq$ 24) 4,318 3,733 326 10 170 199
Complex (key $\leq$ 41) 0 0 0 1 179 733
Extra complex (key $>$ 41) 0 0 0 0 174 976
Average depth of JSON 4.00 3.00 3.48 3.13 5.01 5.19
Average branching factor 1.22 1.44 1.18 1.17 1.41 1.38
Diversity Total # of unique keys 24 12 18 31 275 362
Composite views 0 0 0 0 136 746
Interaction (e.g., zoom, pan) 0 0 0 0 188 1,010
# of chart types 6 6 4 2 10 10
We have collected a new set of 1,981 real-world Vega-Lite specifications. In this section, we present the details of our data collection process.
3.1. Dataset Construction
3.1.1. Search Queries.
3.1.2. Inclusion and Exclusion Criteria.
3.1.3. Post-processing.
3.2. Quantitative Analysis
3.2.1. Benchmarks.
3.2.2. Quality Metrics.
3.2.3. Complexity Levels.
We observe that the existing criteria used to establish the complexity levels of charts are somewhat subjective and may not possess broad applicability (Luo et al., 2021a; Kantharaj et al., 2022; Lundgard and Satyanarayan, 2021). Instead, we suggest using the number of keys as a criterion for categorizing the complexity levels of charts, particularly in the context of Vega-Lite specifications. This is because, as explained above, the number of properties increases proportionately to the number of keys in a specification. To establish the standard number of keys, we refer to the Vega-Lite example gallery dataset (Satyanarayan et al., 2023b) and calculate the quartiles (Q1, Q2, Q3) based on the distribution of the number of keys. These quartiles, specifically 16, 24, and 41, are utilized as reference points to divide the specifications’ level of complexity. For instance, a specification with a total number of keys less than or equal to 16 is classified as ‘simple’ complexity. Likewise, a specification with a total number of keys greater than 16 and less than or equal to 24 is classified as ‘medium’ complexity (Figure 2).
3.2.4. Composite View, Interactivity, and Chart Type Distribution.
3.2.5. Results.
4. VL2NL: NL Generation Framework
Figure 3. LLM Framework to Generate NL Datasets for Visualizations. We start by (b) preprocessing underlying datasets and minifying Vega-Lite specifications. Subsequently, (c) we employ scaffolding and key questions, (e) to generate NL datasets like L1/L2 captions, utterances, and questions. (d) This is followed by score-based paraphrasing, (f) allowing us to produce syntactically paraphrased NL datasets.
Table 3. Prompting techniques to generate each NL dataset. Each prompt is designed by choosing the most appropriate techniques considering their different characteristics.
@@ 4. VL2NL: NL Generation Framework
The goal of our framework is to generate high-quality NL datasets using only Vega-Lite specifications and prompt engineering. VL2NL consists of three stages (Figure 3). First it preprocesses underlying datasets (e.g., csv) and minifies the Vega-Lite specifications. Next, it identifies relevant and accurate information through guided-discovery. Last, it increases syntactic diversity using score-based paraphrasing. To generate each type of NL dataset, we design each prompt to be maximally helpful by selecting the most appropriate strategies (Table 3).
@@ 4.2. Ensuring Accuracy and Relevance
Our framework leverages the concept of guided discovery (Brown and Campione, 1994) based on Chain-of-Thought prompting (Wei et al., 2022) to harness the maximum reasoning capability of LLMs. We employ two strategies of guided discovery: providing scaffolds (Hmelo-Silver et al., 2007). and posing key questions (De Jong and Van Joolingen, 1998). To analyze and integrate the chart semantics necessary for generating a specific NL dataset, we assist LLMs by offering scaffolds. Additionally, we furnish LLMs with key questions to guide their self-directed progress. This maximizes the use of LLMs’ reasoning abilities, allowing them to make decisions on which aspects to focus on and delve into when creating a particular data. Below, we denote each step where we italicize the relevant phrases, and utilize symbols for ($\mathcal{S}$) scaffolding and ($\mathcal{K}$) key question to make them easily identifiable.
@@ 4.2.1. L1 Caption
Considering real-world Vega-Lite specifications, we first understand whether the given chart is ($\mathcal{S}$) a composite view (e.g., layered, trellis, and multiple views), to enable top-down analysis of each chart one by one. The prompt follows a template with three questions to answer: Is it a composite view?; If it is, identify its type among layered, trellis, and multiple views; and determine the number of plots in the chart. Next, it analyzes each chart individually based on the provided scaffold of ($\mathcal{S}$) chart semantics: Data, Transform, Mark, Chart-Type, Encoding, Style, and Interaction, using the information about composite view. Here, we access the underlying dataset to provide ($\mathcal{S}$) fields that are presented in the Vega-Lite specification, along with their synonyms (i.e., ($\mathcal{S}$) titles also found in the same Vega-Lite specification) and their ($\mathcal{S}$) unique values if they are categorical variables. After analyzing all of these semantics, the LLM finally generates the L1 caption by combining them.
@@ 4.2.2. L2 Caption
L2 captions, unlike L1 captions that provide an overall description of the chart, offer the flexibility to selectively focus on specific features that capture the viewer’s interest. To craft informative and insightful captions, we follow a structured approach centered around a key question: ($\mathcal{K}$) What is the most prominent and meaningful feature in the given chart? Once we identify this feature, we delve deeper by exploring the mathematical operations required to analyze it: ($\mathcal{K}$) What is the mathematical operation(s) required to describe the feature? Subsequently, based on these operations, we generate ($\mathcal{K}$) a series of questions to analyze the feature (e.g., for the simple line chart with a red border in Figure 2, the following questions are generated: What was the highest stock price of Google?; What was the lowest stock price of Google?; What is the difference between the highest and lowest stock prices of Google?). This process allows us to create captions that provide valuable insights into the chart’s content. When answering questions, we utilize backing datasets and LangChain (Chase, 2022) to perform required calculation that bolster the factual integrity of the generated captions. This step is crucial, as large language models (LLMs) have been known to produce hallucinations in response to mathematical problems (Ji et al., 2023). Once each question is answered, the collected information is subsequently incorporated into the final prompting stage for generating L2 captions. It’s important to note that, unlike previous work (Tang et al., 2023), we do not use any of the L1 captions when generating L2 captions. Instead, this is performed as an independent process.
@@ 4.2.3. Utterance
Similar to L1 captions, we begin by analyzing whether the chart is a ($\mathcal{S}$) composite view. We then proceed to generate ($\mathcal{S}$) instructions for each plot independently. This process entails creating a comprehensive set of step-by-step instructions for constructing each plot. To enhance readability and user-friendliness, we ensure that each instruction focuses on a single specific action, aligning with the same semantics used when generating L1 captions. For example, in the case of the simple line chart with a red border shown in Figure 2, the following instructions are generated:
@@
However, it is important to note that the generated instructions may sometimes feature overly technical variable names from the chart, which might not align with users’ NL usage patterns. In such cases, we leverage information from synonyms found in the underlying dataset. Specifically, we use ($\mathcal{S}$) title of the Vega-Lite specification and ($\mathcal{S}$) values from the fields to replace the technical terms, resulting in more user-friendly instructions.
Next, we ask a key question to ($\mathcal{K}$) identify primary and secondary information. In this context, we anticipate that LLMs are able to automatically prioritize crucial semantics to paint a comprehensive picture of the chart, such as chart type or encoding, over additional instructions like style or interaction. This thought is based on Wang et al.’s observations (Wang et al., 2022), who noted that the typical workflow for creating visualizations often starts with this information (e.g., ’show me the price over time as a line chart’). Once we have all these components ready, we proceed to generate each type of utterance one by one, adhering to ($\mathcal{S}$) specific rules for each type. For commands, we employ the imperative voice. For queries, we use only variables, fields, attributes, mathematical formulas, abbreviations, and prepositions, while avoiding verbs and articles. For questions, we formulate inquiries in the form of questions. Across all types, we maintain the following rules: express each utterance in a single sentence, utilize only primary information, and keep the language concise and straightforward.